OpenStudy (anonymous):

a camera manufacturer spends $1,800 each day for overhead expenses plus $9 per camera for labor and materials. the cameras sell for $18 each. a) how many cameras must the company sell in one day to equal its daily costs ? b) if the manufacturer can increase production by 50 cameras per day, what would their daily profit be

OpenStudy (anonymous):

(a)Let the number of cameras daily manufactured be N. Cost of cameras sold in 1 day = 18N Total expenses = 1800 + 9N It's given that daily expense should be equal to Cost of cameras sold in 1 day 18N = 1800+9N 9N = 1800 N = 200

OpenStudy (anonymous):

@emadot

OpenStudy (anonymous):

Try (b) yourself :)